Market demand for plastic straws is P = 50 – 0.02Q, and marginal cost MC = 20 + 0.01Q, where Q is a pack of straws and P is price per pack. Each pack of plastic straws further costs the society $15 in plastic pollution. Complete the table below.

  1. The socially optimal quantity of plastic straws is equal to _________________

       b) Briefly explain why in this specific case, a monopoly firm supplying the plastic straws is preferable to a perfectly competitive market supplying the plastic straws use graphs and number , you must provide a clear answer
Answer #1

P = So - 0.02 & Pmc = 20 toolą SmC = 20 +0.01&+ 15 = 35+0.018 Condition for equilibrium SO = 0.029 = 35+ 0.019 15 = 0.039 9 =